# Explicit $7$-torsion in the Tate–Shafarevich groups of genus $2$ Jacobians

### v1

@@ 2.3. The twist $X^{r}_{J[\mathfrak{p}]}(p)$
Note that by Lemma 2.2 we may define the twists $X_{J[\mathfrak{p}]}^{\pm}(p)$ and $X_{J[\bar{\mathfrak{p}}]}^{\pm}(p)$ which (by Lemma 2.1) parametrise elliptic curves which are $(p,\mathfrak{p})$-congruent (respectively $(p,\bar{\mathfrak{p}})$-congruent) to the genus $2$ Jacobian $J/K$.
@@ 2.4. Explicit twisting for $X(7)$
[…]
@@
Composing with the natural $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu$-isomorphism $\Gamma\!_{7}\cong\operatorname{SL}_{2}(\mathbb{Z}/7\mathbb{Z})$ gives a $G_{K}$-equivariant homomorphism $\rho\colon\Gamma\!_{7}\to\operatorname{GL}_{2}(\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu)$. When $\Gamma\!_{7}$ acts on $\mathbb{P}^{2}$ via ${\rho}$ the Klein quartic curve $X(7)\subset\mathbb{P}^{2}$ given by
@@ 3.1. Hilbert’s Theorem 90 is effective
To compute twists of $X(7)$ using Lemma 2.3 we will need to compute matrices which realise a given cocycle as a coboundary. Towards this, note that the standard proof of Hilbert’s Theorem 90 is “nearly” effective. Indeed following [Ser79, Proposition X.3] let $L/K$ be a finite extension of infinite fields, let $\xi\in H^{1}(\operatorname{Gal}(L/K),\operatorname{GL}_{n}(L))$ be a $1$-cocycle, and choose an element $c\in\operatorname{GL}_{n}(L)$. Define a matrix $b\in\mathrm{M}_{n}(L)$ by the Poincaré series
@@
Algorithm 3.1 Generating matrices which realise Hilbert’s Theorem 90.
@@
Remark 3.2.
In step 1 of Algorithm 3.1 the user must choose a matrix $c\in\operatorname{GL}_{n}(L)$. In our application the extension $L/\mathbb{Q}$ will be a finite extension and we will have access to a LLL-reduced $\mathbb{Z}$-basis $\{a_{1},...,a_{m}\}$ for the ring of integers $\mathcal{O}_{L}\subset L$. We choose “small” elements of $L$ by generating a tuple $x_{1},...,x_{m}\in\{0,\pm 1\}$ and considering the element $\sum_{i=1}^{m}x_{i}a_{i}\in\mathcal{O}_{L}$. This approach extends to choosing a matrix $c\in\operatorname{GL}_{n}(L)$ by choosing the $n^{2}$ entries as described (in practice, we also choose almost all $x_{i}$ to be equal to zero).
Remark 3.3.
In principle the recursive step in Algorithm 3.1 may be called many times. In practice, however, we have found very few instances when Algorithm 3.1 fails to terminate in one iteration.
@@ 3.2. The main algorithm
    We compute (a $\mathfrak{q}$-adic approximation to) a matrix $A\in\operatorname{GL}_{3}(L)\subset\operatorname{GL}_{3}(L_{\mathfrak{q}})$ which realises $\xi$ as a coboundary, using Algorithm 3.1.
@@
    We twist $X(7)$ by $A$ and $A^{-T}$ to obtain curves $\mathcal{X}^{\pm}\subset\mathbb{P}^{2}_{L_{\mathfrak{q}}}$. By Lemma 2.3 these twists are $\mathfrak{q}$-adic approximations to the twists $X_{J[\mathfrak{p}]}^{\pm}(7)\subset\mathbb{P}^{2}_{\mathbb{Q}}$. After normalising each equation so that the first non-zero coefficient is equal to $1$, the curves obtained therefore have coefficients in $\mathbb{Q}$ (up to a numerical error).
@@
Let $\alpha$ be a root of $g_{1}(t)$ (so that $L$ is the splitting field of $\mathbb{Q}(\alpha)$). Since the Galois module $J[\mathfrak{p}]$ is irreducible, there exists a subfield $\mathbb{Q}\subset K\subset\mathbb{Q}(\alpha)$ (which is unique up to conjugacy) such that $\mathcal{K}[\mathfrak{p}]$ contains a $\operatorname{Gal}(L/K)$-stable “cyclic subgroup” (or more precisely, the image of a cyclic subgroup of $J[\mathfrak{p}]$). From the properties of the Weil pairing we have $L=\widetilde{K}(\zeta_{7})$, where $\widetilde{K}$ is the Galois closure of $K$ in $L$. We first compute an LLL-reduced basis $\{1,k_{1},...,k_{7}\}$ for $K/\mathbb{Q}$. The elements $k\zeta_{7}^{i}$ span $L/\mathbb{Q}$ where $0\leq i\leq 6$ and $k$ ranges over the $\operatorname{Gal}(L/\mathbb{Q})$-conjugates of $k_{j}$ for each $1\leq j\leq 7$. We then apply Algorithm 3.1 noting Remark 3.2.
@@ 3.3. Computing the polynomials $g_{i}(t)$
[…]
@@ 3.3.1. A numerically stable approach when $D=8$
Fix an isomorphism $\mathcal{O}_{8}\cong\mathbb{Z}[\sqrt{2}]$. The prime number $7$ is a norm from $\mathbb{Z}[\sqrt{2}]$ and we may write $(7)=(3+\sqrt{2})(3-\sqrt{2})$. Let $[\sqrt{2}]\colon J\to J$ denote the multiplication-by-$\sqrt{2}$-map on $J$. The morphism $[\sqrt{2}]$ is a Richelot isogeny and using the approach in [Nic18, Section 5.7] (which is implemented in [Nic]) we determine explicit polynomials giving the morphism $[\sqrt{2}]\colon\mathcal{K}\to\mathcal{K}$ induced by the action of $\sqrt{2}$ on $J$.
Remark 3.5.
@@
Formulae for the multiplication-by-$3$-map $[3]\colon\mathcal{K}\to\mathcal{K}$ are given in [CF96, Section 3.5]. Note that $\mathcal{K}[3+\sqrt{2}]\cup\mathcal{K}[3-\sqrt{2}]$ is exactly the set $\{P\in\mathcal{K}:3P=\sqrt{2}P\}$. By taking successive resultants (and fixing a choice of sign so that $\mathfrak{p}=(3\pm\sqrt{2})$) it is simple to compute polynomials $g_{1}(t),g_{2}(t),g_{3}(t)\in\mathbb{Q}[t]$ satisfying (1).
@@ 3.4. Outputs of the main algorithm
[…]
@@ 4. Proving twists of $X(7)$ are isomorphic to $X_{M}^{\pm}(7)$
For the purpose of proving Theorem 1.1 it suffices to consider only the case when $\mathcal{X}$ has a rational point (i.e., it suffices to recall [Fis16, Lemma 6.2], see Lemma 4.2 below). In Section 4.2 we note how one may prove that a twist is isomorphic to $X_{M}^{\pm}(7)$ more generally.
@@ 4.1. When $\mathcal{X}$ has a rational point
Let $K$ be a field of characteristic zero. Suppose that $\mathcal{X}$ has a $K$-rational point which corresponds (through the moduli interpretation in Lemma 2.4) to an elliptic curve $E/K$ (defined up to quadratic twist) with $j$-invariant $j(E)\neq 0,1728,\infty$. In this case, the following lemma reduces the problem of showing that $\mathcal{X}$ is isomorphic to $X_{M}^{\pm}(7)$ to the problem of showing that $X_{E}(7)$ is isomorphic to $X_{M}^{\pm}(7)$.
Lemma 4.1.
Let $\mathcal{X}/K$ be a twist of $X(7)$ and suppose that there exsists a point $P\in\mathcal{X}(K)$ with $j(P)\neq 0,1728,\infty$. If $E/K$ is an elliptic curve with $j(E)=j(P)$ then $\mathcal{X}$ is isomorphic to $X_{E}(7)$ over $K$.
@@
Lemma 4.2.
Let $K$ be a number field and let $M$ be a $G_{K}$-module which is isomorphic as an abstract group to $(\mathbb{Z}/p\mathbb{Z})^{2}$ for some $p\geq 5$. Suppose that $M$ comes equipped with a ($G_{K}$-equivariant) alternating pairing $M\times M\to\mu_{p}$. Let $E/K$ be an elliptic curve with surjective mod $p$ Galois representation, let $x_{M}\colon M\to M/\{\pm 1\}$, and let $x_{E}\colon E\to\mathbb{P}^{1}$ be the quotient by {±$\{\pm 1\}$. If there exist non-identity elements $P\in M$ and $Q\in E[p]$ such that $K(x_{M}(P))=K(x_{E}(Q))$ then there exists a quadratic twist $E^{d}$ of $E$ such that $M\cong E^{d}[p]$.
@@ 4.2. When $\mathcal{X}$ has no rational points
Proposition 4.4.
Conider any of the data in [Fre, data/twists.m] which consists of
@@
If $\mathcal{X}$ has a $\mathbb{Q}$-rational point of small height corresponding to an elliptic curve $E/\mathbb{Q}$, we apply Lemma 4.2. In each case, applying [Ser72, Proposition 19] at several good primes (or using Zywina’s algorithm [Zyw22]) suffices to show that the mod $7$ Galois representation attached to $E/\mathbb{Q}$ is surjective. In the electronic data we exhibit an explicit isomorphism between the fields $x_{J}(P)$ and $x_{E}(Q)$ for some $P\in J[\mathfrak{p}]$ and $Q\in E[7]$ (note that a minimal polynomial for the extension $\mathbb{Q}(x_{J}(P))/\mathbb{Q}$ was computed in the course of the algorithm in Section 3.3).
@@
The general case proceeds similarly. Taking hyperplane sections of $\mathcal{X}$ we construct non-isomorphic quartic fields $L_{1},L_{2}/\mathbb{Q}$ over which $\mathcal{X}$ obtains a point and such that $L_{i}$ contains no non-trivial subfield for each $i=1,2$ (in particular $L_{1}\cap L_{2}=\mathbb{Q}$ and $L_{1}\cap\mathbb{Q}(\zeta_{7})=L_{2}\cap\mathbb{Q}(\zeta_{7})=\mathbb{Q}$). These points correspond to elliptic curves $E_{1}/L_{1}$ and $E_{2}/L_{2}$ whose mod $7$ Galois representations may be seen to be surjective by applying [Ser72, Proposition 19] at several places of good reduction. Applying Lemma 4.2 as above shows that $E_{1}$ and $E_{2}$ are $(7,\mathfrak{p})$-congruent to $J$, up to a quadratic twist. Since the mod $7$ Galois representations of $E_{1}/L_{1}$ and $E_{2}/L_{2}$ are surjective we have $\mathbb{Q}(\mathcal{K}[\mathfrak{p}])\cap L_{1}L_{2}=\mathbb{Q}$, where $\mathcal{K}=J/\{\pm 1\}$ is the Kummer surface of $J$.
@@
It follows from the construction that $\operatorname{Aut}(X_{J[\mathfrak{p}]}^{r}(7))$ is isomorphic (as a $G_{\mathbb{Q}}$-module) to the group $\operatorname{Aut}_{r}(J[\mathfrak{p}])/\{\pm 1\}$ consisting of automorphisms of $J[\mathfrak{p}]$ which are symplectic with respect to $(e_{J,7})^{r}$. Therefore, the field of definition of the automorphisms of $X_{J[\mathfrak{p}]}^{\pm}(7)$ is equal to $\mathbb{Q}(\mathcal{K}[\mathfrak{p}])$ and $X_{J[\mathfrak{p}]}^{\pm}(7)$ admits no non-trivial automorphisms over $L_{1}L_{2}$ that are not defined over $\mathbb{Q}$. Suppose there is such an automorphism $\tau$ defined over $\mathbb{Q}$. Since the mod $7$ Galois representation attached to $J[\mathfrak{p}]$ is surjective ($J[\mathfrak{p}]$ is isomorphic over $L_{i}$ to a quadratic twist of $E_{i}[7]$ for each $i=1,2$) the element $\tau$ is contained in the centre of $\operatorname{Aut}_{r}(J[\mathfrak{p}])/\{\pm 1\}$ which is isomorphic to $\operatorname{PSL}_{2}(\mathbb{Z}/7\mathbb{Z})$ as an abstract group. Therefore $\tau$ is the identity and the claim follows from Lemma 4.3. ∎
5. Proving $(7,\mathfrak{p})$-congruences and Theorem 1.1

### v2

@@ 2.3. The twist $X^{r}_{J[\mathfrak{p}]}(p)$
Note that by Lemma 8 we may define the twists $X_{J[\mathfrak{p}]}^{\pm}(p)$ and $X_{J[\bar{\mathfrak{p}}]}^{\pm}(p)$ which (by Lemma 7) parametrise elliptic curves which are $(p,\mathfrak{p})$-congruent (respectively $(p,\bar{\mathfrak{p}})$-congruent) to the genus $2$ Jacobian $J/K$.
@@ 2.4. Explicit twisting for $X(7)$
Following [Fis14] we give an explicit description for the twists $X_{M}^{\pm}(7)$ as plane quartic curves. Recall that we write $\Gamma\!_{7}$ for the automorphism group (scheme) of the symplectic abelian group $\mu_{7}\times\mathbb{Z}/7\mathbb{Z}$. Following Klein [Kle78, Elk99, Fis14] consider the representation $\operatorname{SL}_{2}(\mathbb{Z}/7\mathbb{Z})\to\operatorname{GL}_{3}(\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu)$ which maps the generators $S=\left(\begin{smallmatrix}0&-1\\ 1&0\end{smallmatrix}\right)$ and $T=\left(\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right)$ to
@@
Composing with the natural $\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu$-isomorphism $\Gamma\!_{7}\cong\operatorname{SL}_{2}(\mathbb{Z}/7\mathbb{Z})$ gives a $G_{K}$-equivariant homomorphism $\rho\colon\Gamma\!_{7}\to\operatorname{GL}_{3}(\mkern 1.5mu\overline{\mkern-1.5muK\mkern-1.5mu}\mkern 1.5mu)$. When $\Gamma\!_{7}$ acts on $\mathbb{P}^{2}$ via $\rho$} the Klein quartic curve $X(7)\subset\mathbb{P}^{2}$ given by
@@ 3.1. Hilbert’s Theorem 90 is effective
To compute twists of $X(7)$ using Lemma 9 we will need to compute matrices which realise a given cocycle as a coboundary. Towards this, note that the standard proof of Hilbert’s Theorem 90 is “nearly” effective. Indeed following [Ser79, Proposition X.3] let $L/K$ be a finite extension of infinite fields, let $\xi\in H^{1}(\operatorname{Gal}(L/K),\operatorname{GL}_{n}(L))$ be a $1$-cocycle, and choose an element $c\in\operatorname{GL}_{n}(L)$. Define a matrix $b\in\mathrm{M}_{n}(L)$ by the Poincaré series
@@
Algorithm 11 Generating matrices which realise Hilbert’s Theorem 90.
@@
Remark 12.
In step 1 of Algorithm 11 the user must choose a matrix $c\in\operatorname{GL}_{n}(L)$. In our application the extension $L/\mathbb{Q}$ will be a finite extension and we will have access to a LLL-reduced $\mathbb{Z}$-basis $\{a_{1},...,a_{m}\}$ for the ring of integers $\mathcal{O}_{L}\subset L$. We choose “small” elements of $L$ by generating a tuple $x_{1},...,x_{m}\in\{0,\pm 1\}$ and considering the element $\sum_{i=1}^{m}x_{i}a_{i}\in\mathcal{O}_{L}$. This approach extends to choosing a matrix $c\in\operatorname{GL}_{n}(L)$ by choosing the $n^{2}$ entries as described (in practice, we also choose almost all $x_{i}$ to be equal to zero).
Remark 13.
In principle the iteration in Step 6 in Algorithm 11 may be called many times. In practice, however, we have found very few instances when Algorithm 11 fails to terminate in one iteration.
@@ 3.2. The main algorithm
    We compute (a $\mathfrak{q}$-adic approximation to) a matrix $A\in\operatorname{GL}_{3}(L)\subset\operatorname{GL}_{3}(L_{\mathfrak{q}})$ which realises $\xi$ as a coboundary, using Algorithm 11.
@@
    We twist $X(7)$ by $A$ and $A^{-T}$ to obtain curves $\mathcal{X}^{\pm}\subset\mathbb{P}^{2}_{L_{\mathfrak{q}}}$. By Lemma 9 these twists are $\mathfrak{q}$-adic approximations to the twists $X_{J[\mathfrak{p}]}^{\pm}(7)\subset\mathbb{P}^{2}_{\mathbb{Q}}$. After normalising each equation so that the first non-zero coefficient is equal to $1$, the curves obtained therefore have coefficients in $\mathbb{Q}$ (up to a numerical error).
@@
Let $\alpha$ be a root of $g_{1}(t)$ (so that $L$ is the splitting field of $\mathbb{Q}(\alpha)$). Since the Galois module $J[\mathfrak{p}]$ is irreducible, there exists a subfield $\mathbb{Q}\subset K\subset\mathbb{Q}(\alpha)$ (which is unique up to conjugacy) such that $\mathcal{K}[\mathfrak{p}]$ contains a $\operatorname{Gal}(L/K)$-stable “cyclic subgroup” (or more precisely, the image of a cyclic subgroup of $J[\mathfrak{p}]$). From the properties of the Weil pairing we have $L=\widetilde{K}(\zeta_{7})$, where $\widetilde{K}$ is the Galois closure of $K$ in $L$. We first compute an LLL-reduced basis $\{1,k_{1},...,k_{7}\}$ for $K/\mathbb{Q}$. The elements $k\zeta_{7}^{i}$ span $L/\mathbb{Q}$ where $0\leq i\leq 6$ and $k$ ranges over the $\operatorname{Gal}(L/\mathbb{Q})$-conjugates of $k_{j}$ for each $1\leq j\leq 7$. We then apply Algorithm 11 noting Remark 12.
@@ 3.3. Computing the polynomials $g_{i}(t)$
Our approach follows that of Fisher [Fis16, Theorem 6.3]. Using the analytic Jacobian machinery in Magma (in particular the functions AnalyticJacobian and EndomorphismRing) we compute complex approximations to a $\mathfrak{p}$-torsion divisor $\mathfrak{D}=(x_{1},y_{1})+(x_{2},y_{2})-(\infty^{+}+\infty^{-})\in J(\mkern 1.5mu\overline{\mkern-1.5mu\mathbb{Q}\mkern-1.5mu}\mkern 1.5mu)$.
@@ 3.3.1. A numerically stable approach when $D=8$
Fix an isomorphism $\mathcal{O}_{8}\cong\mathbb{Z}[\sqrt{2}]$. The prime number $7$ is a norm from $\mathbb{Z}[\sqrt{2}]$ and we may write $(7)=(3+\sqrt{2})(3-\sqrt{2})$. Let $\sqrt{2}]\colon J\to J$ denote the multiplication-by-$\sqrt{2}$-map on $J$. The morphism $\sqrt{2}]$ is a Richelot isogeny and using the approach in [Nic18, Section 5.7] (which is implemented in [Nic]) we determine explicit polynomials giving the morphism $\sqrt{2}]\colon\mathcal{K}\to\mathcal{K}$ induced by the action of $\sqrt{2}$ on $J$.
Remark 15.
@@
Formulae for the multiplication-by-$3$-map $3]\colon\mathcal{K}\to\mathcal{K}$ are given in [CF96, Section 3.5]. Note that $\mathcal{K}[3+\sqrt{2}]\cup\mathcal{K}[3-\sqrt{2}]$ is exactly the set $\{P\in\mathcal{K}:3P=\sqrt{2}P\}$. By taking successive resultants (and fixing a choice of sign so that $\mathfrak{p}=(3\pm\sqrt{2})$) it is simple to compute polynomials $g_{1}(t),g_{2}(t),g_{3}(t)\in\mathbb{Q}[t]$ satisfying (1).
@@ 3.4. Outputs of the main algorithm
A generic family of genus $2$ curves $C/\mathbb{Q}$ whose Jacobians have RM by $\mathcal{O}_{8}$ are given by Bending [Ben98, Ben99], who also records many examples of small conductor in [Ben98, Appendix A]. Bending’s family is given by a triple of parameters $A,P,Q\in\mathbb{Q}$. It is simple to search for further examples of small conductor (noting from [Ben98, Section 6.3] that it is often useful to specialise at $P\in\{\pm 1,\pm 1/2,\pm 1/3,\pm 1/5\}$). Combining these with the examples found in the LMFDB [LMFDB] we obtain a small (non-exhaustive) database of curves with RM by $\mathbb{Z}[\sqrt{2}]$ and whose Jacobians have conductor $\sqrt{N_{J}}\leq 500\,000$ (these may be found in [Fre]).
Similar generic families are provided for each $D=8$, $29$, $37$, $44$, and $53$ in [CFM24] building on work of Elkies–Kumar [EK14], who compute the moduli of such curves for all fundamental discriminants $D<100$. Some examples of curves with RM by $\mathcal{O}_{D}$ and with small conductor are recorded in [EK14]. We record a (non-exhaustive) list of such curves with $\sqrt{N_{J}}\leq 500\,000$ in [Fre]. Note that when $D>17$ the moduli space of curves with RM by $\mathcal{O}_{D}$ is not rational, so examples are sparser than when $D=8$.
@@ 4. Proving twists of $X(7)$ are isomorphic to $X_{M}^{\pm}(7)$
For the purpose of proving Theorem 1 it suffices to consider only the case when $\mathcal{X}$ has a rational point (i.e., it suffices to recall [Fis16, Lemma 6.2], see Lemma 18 below). In Section 4.2 we note how one may prove that a twist is isomorphic to $X_{M}^{\pm}(7)$ more generally.
@@ 4.1. When $\mathcal{X}$ has a rational point
Let $K$ be a field of characteristic zero. Suppose that $\mathcal{X}$ has a $K$-rational point which corresponds (through the moduli interpretation in Lemma 10) to an elliptic curve $E/K$ (defined up to quadratic twist) with $j$-invariant $j(E)\neq 0,1728,\infty$. In this case, the following lemma reduces the problem of showing that $\mathcal{X}$ is isomorphic to $X_{M}^{\pm}(7)$ to the problem of showing that $X_{E}(7)$ is isomorphic to $X_{M}^{\pm}(7)$.
Lemma 17.
Let $\mathcal{X}/K$ be a twist of $X(7)$ and suppose that there exists a point $P\in\mathcal{X}(K)$ with $j(P)\neq 0,1728,\infty$. If $E/K$ is an elliptic curve with $j(E)=j(P)$ then $\mathcal{X}$ is isomorphic to $X_{E}(7)$ over $K$.
@@
Lemma 18.
Let $K$ be a number field and let $M$ be a $G_{K}$-module which is isomorphic as an abstract group to $(\mathbb{Z}/p\mathbb{Z})^{2}$ for some $p\geq 5$. Suppose that $M$ comes equipped with a ($G_{K}$-equivariant) alternating pairing $M\times M\to\mu_{p}$. Let $E/K$ be an elliptic curve with surjective mod $p$ Galois representation, let $x_{M}\colon M\to M/\{\pm 1\}$, and let $x_{E}\colon E\to\mathbb{P}^{1}$ be the quotient by $\{\pm 1\}$. If there exist non-identity elements $P\in M$ and $Q\in E[p]$ such that $K(x_{M}(P))=K(x_{E}(Q))$ then there exists a quadratic twist $E^{d}$ of $E$ such that $M\cong E^{d}[p]$.
@@ 4.2. When $\mathcal{X}$ has no rational points
Proposition 20.
Consider any of the data in [Fre, data/twists.m] which consists of
@@
If $\mathcal{X}$ has a $\mathbb{Q}$-rational point of small height corresponding to an elliptic curve $E/\mathbb{Q}$, we apply Lemma 18. In each case, applying [Ser72, Proposition 19] at several good primes (or using Zywina’s algorithm [Zyw22]) suffices to show that the mod $7$ Galois representation attached to $E/\mathbb{Q}$ is surjective. In the electronic data we exhibit an explicit isomorphism between the fields $x_{J}(P)$ and $x_{E}(Q)$ for some $P\in J[\mathfrak{p}]$ and $Q\in E[7]$ (note that a minimal polynomial for the extension $\mathbb{Q}(x_{J}(P))/\mathbb{Q}$ was computed in the course of the algorithm in Section 3.3).
@@
The general case proceeds similarly. Taking hyperplane sections of $\mathcal{X}$ we construct non-isomorphic quartic fields $L_{1},L_{2}/\mathbb{Q}$ over which $\mathcal{X}$ obtains a point and such that $L_{i}$ contains no non-trivial subfield for each $i=1,2$ (in particular $L_{1}\cap L_{2}=\mathbb{Q}$ and $L_{1}\cap\mathbb{Q}(\zeta_{7})=L_{2}\cap\mathbb{Q}(\zeta_{7})=\mathbb{Q}$). These points correspond to elliptic curves $E_{1}/L_{1}$ and $E_{2}/L_{2}$ whose mod $7$ Galois representations may be seen to be surjective by applying [Ser72, Proposition 19] at several places of good reduction. Applying Lemma 18 as above shows that $E_{1}$ and $E_{2}$ are $(7,\mathfrak{p})$-congruent to $J$, up to a quadratic twist. Since the mod $7$ Galois representations of $E_{1}/L_{1}$ and $E_{2}/L_{2}$ are surjective we have $\mathbb{Q}(\mathcal{K}[\mathfrak{p}])\cap L_{1}L_{2}=\mathbb{Q}$, where $\mathcal{K}=J/\{\pm 1\}$ is the Kummer surface of $J$.
@@
It follows from the construction that $\operatorname{Aut}(X_{J[\mathfrak{p}]}^{r}(7))$ is isomorphic (as a $G_{\mathbb{Q}}$-module) to the group $\operatorname{Aut}_{r}(J[\mathfrak{p}])/\{\pm 1\}$ consisting of automorphisms of $J[\mathfrak{p}]$ which are symplectic with respect to $(e_{J,7})^{r}$. Therefore, the field of definition of the automorphisms of $X_{J[\mathfrak{p}]}^{\pm}(7)$ is equal to $\mathbb{Q}(\mathcal{K}[\mathfrak{p}])$ and $X_{J[\mathfrak{p}]}^{\pm}(7)$ admits no non-trivial automorphisms over $L_{1}L_{2}$ that are not defined over $\mathbb{Q}$. Suppose there is such an automorphism $\tau$ defined over $\mathbb{Q}$. Since the mod $7$ Galois representation attached to $J[\mathfrak{p}]$ is surjective ($J[\mathfrak{p}]$ is isomorphic over $L_{i}$ to a quadratic twist of $E_{i}[7]$ for each $i=1,2$) the element $\tau$ is contained in the centre of $\operatorname{Aut}_{r}(J[\mathfrak{p}])/\{\pm 1\}$ which is isomorphic to $\operatorname{PSL}_{2}(\mathbb{Z}/7\mathbb{Z})$ as an abstract group. Therefore $\tau$ is the identity and the claim follows from Lemma 19. ∎
5. Proving $(7,\mathfrak{p})$-congruences and Theorem 1